Barry has three sisters. The average age of the three sisters is 27. The average age of Barry and his three sisters is 28. What is Barry's age?
Answer by lwsshak3(11628)

Barry has three sisters. The average age of the three sisters is 27. The average age of Barry and his three sisters is 28. What is Barry's age?
***
let x=Barry's age
(3*27+x)/4=28
81+x=112
x=112-81=31
What is Barry's age? 31